OpenStudy (anonymous):

Calculate the area of triangle WXY with altitude YZ, given W(2, −1), X(6, 3), Y(7, 0), and Z(5, 2). 8 square units 9.4 square units 7.7 square units 12 square units

OpenStudy (amistre64):

you need to find the distances for a base and height base is WX and height is YZ what is our distance formula?

OpenStudy (amistre64):

hmm i wonder if we can generalize it so we dont have to sqrt A = 1/2 bh if we find b^2 and h^2 A' = 1/2 (bh)^2 either way theres a sqrt but waitning till the end might be useful at times

OpenStudy (anonymous):

When i drew the picture it didn't make sense @amistre64

OpenStudy (amistre64):

if pictures, we can make a rectangle about the triangle 

OpenStudy (amistre64):

move it to the origin add -2,1 to it all